OpenStudy (anonymous):

Find the solutions of g(x)=x^2+6x+1

OpenStudy (anonymous):

explain each step.

OpenStudy (zzr0ck3r):

I assume you mean the zeros of g(x)?

OpenStudy (zzr0ck3r):

or the solution to \(x^2+6x+1=0\)

OpenStudy (anonymous):

yes but idk what to do from that poing @zzr0ck3r

OpenStudy (zzr0ck3r):

have you seen this \(\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) before ?

OpenStudy (anonymous):

yes but I never really got what it was used for

OpenStudy (zzr0ck3r):

its used for this question

OpenStudy (anonymous):

oh okay

OpenStudy (zzr0ck3r):

it will give the zeros of the polynomial \(ax^2+bx+c\) for you \(a=1, b=6,\) and \(c=1\)

OpenStudy (anonymous):

so then you plug the number in and whatever you get is the answer?

OpenStudy (zzr0ck3r):

yeah

OpenStudy (zzr0ck3r):

\(\frac{-6\pm\sqrt{6^2-4*1*1}}{2*1}\)

OpenStudy (zzr0ck3r):

you will get two answers

OpenStudy (anonymous):

how?

OpenStudy (zzr0ck3r):

the \(\pm\) will lead to two answers

OpenStudy (zzr0ck3r):

what that reall is, is \(\frac{-6+\sqrt{6^2-4*1*1}}{2*1}\) and \(\frac{-6-\sqrt{6^2-4*1*1}}{2*1}\)

OpenStudy (zzr0ck3r):

really*

OpenStudy (anonymous):

ooooh okay. so how would you solve this part \[-6-\sqrt{6^{2}-4*1*1}\] I don't do good with radicals

OpenStudy (zzr0ck3r):

what is 6^2?

OpenStudy (anonymous):

6 squared

OpenStudy (zzr0ck3r):

\(\frac{-6+\sqrt{6^2-4*1*1}}{2*1}=\frac{-6+\sqrt{36-4}}{2}=\frac{-6+\sqrt{32}}{2*1}=\frac{-6+\sqrt{16*2}}{2}=\\\frac{-6+\sqrt{16}\sqrt{2}}{2}=\frac{-6+4\sqrt{2}}{2}=-3+2\sqrt{2}\)

OpenStudy (anonymous):

its the same thing with the other equation too?